A reduction in concentration of A from 120 mg/L to 2
mg/L in two hours. It is a zero order reaction determine the value of
the rate constant with appropriate units. Repeat the analysis if the reaction is known to follow first-order kinetics.

Show all work to include conversions. Answer will be rated after it has been verified correct.
Answer #1

Initial concentration = 120 mg/L

Final concentration = 2 mg/L

Time = 2 hr

For the zero order reaction :-

A= \left [ A_{0} \right ]-k.t

putting the values we get

2 mg/L= 120mg/L-k\times 2 hr

k= 59mg/L.hr

For the first order reaction the formula is:-

ln\left [ A \right ]=-k.t +ln\left [ A_{0} \right ]

putting values we get:-

ln\left [ 2\right ]=-k\times 2 hr +ln\left [ 120 \right ]

0.693=-k\times 2 hr +4.787

k= 2.047 hr^{-1}